# package/before/amsmath

### 14 Dec 2023 v1

@@ 6 Conclusion
In this paper, we extended the optimal trade execution problem under the Markowitz criteria to a time-consistent optimal trade execution problem for multiple assets and multiple agents setting. This is done by formulating the original mean-variance problem into its dual representation coupled with an auxiliary HJB equation Aivaliotis and Veretennikov [2018]. To extend this into higher dimensions, we reformulate the auxiliary HJB equation into a BSDE that can be solved using a neural network. We present a deep residual self-attention network to solve the BSDE presented in (21). This allows us to use a deep residual network that learns temporal dependencies. This overcomes the issue of saving weights for every timestep $N$ and overcomes the vanishing gradient problem of recurrent networks Na and Wan [2023].
@@
In this paper, we attempted to cover as much as we can. However, there are many different aspects left to explore. The model we used can be updated to include models for limit order books and trade volumes Alfonsi et al. [2010]; Cartea and Jaimungal [2016].

### 19 Mar 2024 v2

@@ 6 Conclusion
In this paper, we extended the optimal trade execution problem under the Markowitz criteria to a time-consistent optimal trade execution problem for multiple assets and multiple agents setting. This is done by formulating the original mean-variance problem into its dual representation coupled with an auxiliary HJB equation [Aivaliotis and Veretennikov 2018]. To extend this into higher dimensions, we reformulate the auxiliary HJB equation into a BSDE that can be solved using a neural network. We present a deep residual self-attention network to solve the BSDE presented in (19). This allows us to use a deep residual network that learns temporal dependencies. This overcomes the issue of saving weights for every timestep $N$ and overcomes the vanishing gradient problem of recurrent networks [Na and Wan 2023].
@@
In this paper, we attempted to cover as much as we can. However, there are many different aspects left to explore. The model we used can be updated to include models for limit order books and trade volumes [Alfonsi et al. 2010; Cartea and Jaimungal 2016].